OpenStudy (anonymous):

Find dy/dx of log (1+√x)^2

OpenStudy (sriram):

y=2 [log1*(1/2)logx] dy=log1*1/x *dx dy/dx=0

OpenStudy (anonymous):

log (1+√x)^2 = ln (1+√x)^2/ln 10 We'll differentiate with respect to x: [ln (1+√x)^2/ln 10]' = (1/ln 10)*(1/(1+√x)^2)*[(1+√x)^2]' [ln (1+√x)^2/ln 10]' = 2*(1+√x)*(1+√x)' /(ln 10)*(1+√x)^2 [ln (1+√x)^2/ln 10]' = 2*(1+√x)/2√x*(ln 10)*(1+√x)^2 dy/dx = [ln (1+√x)^2/ln 10]' =(1+√x) /√x*(ln 10)*(1+√x)^2